My team of basketball has won 40 percent of our games from 15 games. How many more games in a row do we play to have a 55 percent average for wins?
Answer by stanbon(75887)

My team of basketball has won 40 percent of our games from 15 games. How many more games in a row do we play to have a 55 percent average for wins?
------------
Let x be that number of games.
-----
Equation:
(0.4*15 + x) = 0.55*(40+x)
-----
(6+x) = 22 + 0.55x
0.45 = 16
x = 35.56
---
Answer 36 games when rounded up.
================
Cheers,
Stan H.